How is this done?
Solve the following absolute value. Remember there will be two answers.
Show all work. 6 - |4x - 3| = -19

Answer by Earlsdon(6294)

First, isolate the absolute-value part:
6-|4x-3| = -19 Subtract 6 from both sides.
-|4x-3| = -25 Multiply both sides by -1
|4x-3| = 25 Now remove the absolute-value bars to get the two equations.
1) 4x-3 = 25
2) 4x-3 = -25
1) 4x = 28 so x = 7
2) 4x = -22 so x = -5.5